# Full author list in Contributions

### v1

@@ 1 Introduction
Reinforcement learning from human feedback (RLHF) has become the dominant paradigm for aligning language models with human values [9, 23, 3]. Reward models trained via RLHF achieve 95% accuracy on RewardBench’s objective tasks [14], which emphasize safety violations, factual errors, and instruction-following. However, our benchmark reveals a critical limitation: when we systematically remove objective quality signals (grammatical errors, factual mistakes, length differences), sequence-based reward models—the dominant architecture in production RLHF systems—collapse to 52.7% accuracy on writing preference tasks, barely above random chance. This 42-percentage-point degradation indicates that current preference learning primarily optimizes for error detection rather than recognition of subjective creative quality—a fundamental misalignment between training objectives and the aesthetic judgment required for creative tasks.
However, writing tasks constitute over 40% of language model interactions [22, 2], spanning creative fiction, persuasive essays, and personal expression where subjective quality matters more than objective correctness. Yet our evaluation infrastructure remains anchored in verifiable metrics. RewardBench [14] conflates safety with preference; WritingBench mixes creative with functional tasks [31]; LitBench uses Reddit upvotes as quality proxies [11]. Moreover, existing benchmarks predominantly focus on English, leaving cross-lingual preference evaluation, particularly for languages with distinct rhetorical traditions like Chinese, largely unexplored. Recent theoretical work warns of “reward hacking” where models exploit spurious correlations rather than learning genuine preferences [24].
@@
Our evaluation of 21 models, comprising 7 reward models and 14 language models as judges spanning open-source and proprietary families, reveals fundamental limitations in current preference learning approaches. Sequence-based reward models, the dominant architecture in production RLHF systems [25], achieve only 52.7% mean accuracy across both languages, while zero-shot language model judges [32] reach 53.9%. Both results are statistically indistinguishable from random chance. In stark contrast, generative reward models that produce explicit reasoning chains [6] achieve 81.8% accuracy. This 29-percentage-point gap indicates that subjective preference modeling requires structured intermediate reasoning rather than direct pattern matching. Moreover, all architectures exhibit severe genre instability (mean standard deviation 10.9%), with individual models ranging from 18.2% to 81.8% accuracy across categories, suggesting reliance on superficial heuristics rather than generalizable aesthetic principles. Notably, these failures persist across model scales: 27B parameter models show no consistent improvement over 8B variants, and reasoning-enhanced LLMs (Claude-4-Opus-thinking, OpenAI-o3) provide no advantage over standard architectures.
@@ 3.2 Models
  • Sequence Classifiers: Nvidia/AceMath-7B-RM [19], RM-Mistral-7B [10], Skywork-Reward-Llama-3.1-8B [17], Skywork-Reward-Gemma-2-27B [17]
  • Generative RMs: RM-R1-DeepSeek-Qwen-7B, RM-R1-DeepSeek-Qwen-14B, RM-R1-Qwen2.5-7B [6, 30, 13]
Language Model Judges. We evaluate 14 models including reasoning-enhanced variants (Claude-4-{Opus, Sonnet}-thinking [1], Doubao-1.6-thinking [26], OpenAI-o3-high [21]) and standard models (Gemini-2.5-{Flash, Pro} [29], DeepSeek-R1 [13], ByteDance-Seed-1.6 [28], Doubao-{1.5-Lite, 1.5-Pro, 1.6-flash} [5], Qwen-3-235B [30], OpenAI-o4-mini).
@@ 4.1 Reward Model Performance
We evaluate seven reward models across WritingPreferenceBench. Following RewardBench [14], models divide into sequence classifiers (discriminative heads on language models) and generative reward models. However, our results reveal that the RM-R1 series [6] represents a distinct category—generative models that produce reasoning chains before preference judgments, diverging from both traditional classifiers and DPO-based approaches evaluated in RewardBench. Table 1 presents accuracy across eight writing genres in English and Chinese.
@@ 4.2 Language Model Judge Performance
Implications for LLM-as-judge paradigm. The systematic underperformance of LLM judges relative to specialized reward models challenges the widespread adoption of LLM-as-judge evaluation [32]. Mean accuracy of 53.9%—barely above random—indicates that zero-shot prompting cannot elicit reliable preference judgments for subjective tasks. The failure of reasoning-enhanced models further suggests that the limitation is not computational but representational: without explicit preference training, even advanced LLMs default to surface-level heuristics rather than genuine quality assessment.
@@ 6 Related Work
Preference Learning and Evaluation Benchmarks. Modern preference learning originated with Christiano et al. [9], scaled through InstructGPT [23] and RLHF [9]. Subsequent benchmarks [12, 3, 25, 7] and comprehensive evaluations like RewardBench [14], AlpacaEval [16], and MT-Bench [32] measure conversation, reasoning, and instruction-following. However, these benchmarks conflate objective correctness with subjective preference. RewardBench achieves 95% on safety but cannot evaluate aesthetic judgment; MT-Bench measures factual accuracy, not creative quality. Our work reveals that models excelling on these benchmarks fail catastrophically (52.7% accuracy) when objective signals are neutralized.
Evaluating Creative and Subjective Writing. Creative writing evaluation faces inherent subjectivity challenges [8, 4, 20, 15, 27]. Early reference-based metrics (BLEU, ROUGE) fail for open-ended generation. Recent benchmarks make progress but retain critical limitations. LitBench [11] uses Reddit upvotes—confounding preference with popularity and timing—and covers only English. WritingBench [31] spans six domains but mixes subjective creative tasks with objective functional ones (Academic & Engineering, Finance & Business). AlignBench [18] evaluates Chinese LLM alignment but focuses on general capabilities rather than creative preference. WritingPreferenceBench advances beyond these by: (1) systematically neutralizing objective confounds through human validation, (2) providing cross-lingual coverage with consistent methodology, and (3) isolating purely subjective quality discrimination where prior benchmarks conflate multiple signals.
@@ 7 Conclusion
These results have theoretical and practical implications for preference learning. The 30 percentage point performance gap between architectures challenges the direct preference optimization paradigm [25] and suggests that subjective domains require fundamentally different inductive biases than objective tasks. The failure of scale to improve performance (27B models underperform 7B variants) and the inability of reasoning-enhanced LLMs to surpass task-specific training indicate that current scaling laws may not apply to subjective preference modeling. Future work should investigate hybrid architectures combining the computational efficiency of discriminative models with the representational capacity of generative reasoning, and develop training objectives that explicitly encourage genre-invariant preference learning.

### v2

@@ 1 Introduction
Reinforcement learning from human feedback (RLHF) has become the dominant paradigm for aligning language models with human values [christiano2017deep, ouyang2022training, bai2022constitutional]. Reward models trained via RLHF achieve 95% accuracy on RewardBench’s objective tasks [lambert2024rewardbench], which emphasize safety violations, factual errors, and instruction-following. However, our benchmark reveals a critical limitation: when we systematically remove objective quality signals (grammatical errors, factual mistakes, length differences), sequence-based reward models—the dominant architecture in production RLHF systems—collapse to 52.7% accuracy on writing preference tasks, barely above random chance. This 42-percentage-point degradation indicates that current preference learning primarily optimizes for error detection rather than recognition of subjective creative quality—a fundamental misalignment between training objectives and the aesthetic judgment required for creative tasks.
However, writing tasks constitute over 40% of language model interactions [openai2025usage, anthropic2025economic], spanning creative fiction, persuasive essays, and personal expression where subjective quality matters more than objective correctness. Yet our evaluation infrastructure remains anchored in verifiable metrics. RewardBench [lambert2024rewardbench] conflates safety with preference; WritingBench mixes creative with functional tasks [wu2025writingbench]; LitBench uses Reddit upvotes as quality proxies [fein2025litbenchbenchmarkdatasetreliable]. Moreover, existing benchmarks predominantly focus on English, leaving cross-lingual preference evaluation, particularly for languages with distinct rhetorical traditions like Chinese, largely unexplored. Recent theoretical work warns of “reward hacking” where models exploit spurious correlations rather than learning genuine preferences [pan2022effects].
@@
Our evaluation of 21 models, comprising 7 reward models and 14 language models as judges spanning open-source and proprietary families, reveals fundamental limitations in current preference learning approaches. Sequence-based reward models, the dominant architecture in production RLHF systems [rafailov2023direct], achieve only 52.7% mean accuracy across both languages, while zero-shot language model judges [zheng2023judging] reach 53.9%. Both results are statistically indistinguishable from random chance. In stark contrast, generative reward models that produce explicit reasoning chains [chen2025rm] achieve 81.8% accuracy. This 29-percentage-point gap indicates that subjective preference modeling requires structured intermediate reasoning rather than direct pattern matching. Moreover, all architectures exhibit severe genre instability (mean standard deviation 10.9%), with individual models ranging from 18.2% to 81.8% accuracy across categories, suggesting reliance on superficial heuristics rather than generalizable aesthetic principles. Notably, these failures persist across model scales: 27B parameter models show no consistent improvement over 8B variants, and reasoning-enhanced LLMs (Claude-4-Opus-thinking, OpenAI-o3) provide no advantage over standard architectures.
@@ 3.2 Models
  • Sequence Classifiers: Nvidia/AceMath-7B-RM [acemath2024], RM-Mistral-7B [dong2023raft], Skywork-Reward-Llama-3.1-8B [liu2024skywork], Skywork-Reward-Gemma-2-27B [liu2024skywork]
  • Generative RMs: RM-R1-DeepSeek-Qwen-7B, RM-R1-DeepSeek-Qwen-14B, RM-R1-Qwen2.5-7B [chen2025rm, qwen2.5, guo2025deepseek]
Language Model Judges. We evaluate 14 models including reasoning-enhanced variants (Claude-4-{Opus, Sonnet}-thinking [anthropic2025claude4], Doubao-1.6-thinking [seed2025seed1], OpenAI-o3-high [openai2025o3high]) and standard models (Gemini-2.5-{Flash, Pro} [geminiteam2024gemini15unlockingmultimodal], DeepSeek-R1 [guo2025deepseek], ByteDance-Seed-1.6 [bytedance2025seed1.6], Doubao-{1.5-Lite, 1.5-Pro, 1.6-flash} [bytedance_doubao15_pro_2025], Qwen-3-235B [qwen2.5], OpenAI-o4-mini).
@@ 4.1 Reward Model Performance
We evaluate seven reward models across WritingPreferenceBench. Following RewardBench [lambert2024rewardbench], models divide into sequence classifiers (discriminative heads on language models) and generative reward models. However, our results reveal that the RM-R1 series [chen2025rm] represents a distinct category—generative models that produce reasoning chains before preference judgments, diverging from both traditional classifiers and DPO-based approaches evaluated in RewardBench. Table 1 presents accuracy across eight writing genres in English and Chinese.
@@ 4.2 Language Model Judge Performance
Implications for LLM-as-judge paradigm. The systematic underperformance of LLM judges relative to specialized reward models challenges the widespread adoption of LLM-as-judge evaluation [zheng2023judging]. Mean accuracy of 53.9%—barely above random—indicates that zero-shot prompting cannot elicit reliable preference judgments for subjective tasks. The failure of reasoning-enhanced models further suggests that the limitation is not computational but representational: without explicit preference training, even advanced LLMs default to surface-level heuristics rather than genuine quality assessment.
@@ 6 Related Work
Preference Learning and Evaluation Benchmarks. Modern preference learning originated with Christiano et al. [christiano2017deep], scaled through InstructGPT [ouyang2022training] and RLHF [christiano2017deep]. Subsequent benchmarks [gao2023scaling, bai2022constitutional, rafailov2023direct, chiang2024chatbot] and comprehensive evaluations like RewardBench [lambert2024rewardbench], AlpacaEval [alpaca_eval], and MT-Bench [zheng2023judging] measure conversation, reasoning, and instruction-following. However, these benchmarks conflate objective correctness with subjective preference. RewardBench achieves 95% on safety but cannot evaluate aesthetic judgment; MT-Bench measures factual accuracy, not creative quality. Our work reveals that models excelling on these benchmarks fail catastrophically (52.7% accuracy) when objective signals are neutralized.
Evaluating Creative and Subjective Writing. Creative writing evaluation faces inherent subjectivity challenges [chodorow2007detection, burstein2003finding, miltsakaki2000automated, li2018coherence, tay2018skipflow]. Early reference-based metrics (BLEU, ROUGE) fail for open-ended generation. Recent benchmarks make progress but retain critical limitations. LitBench [fein2025litbenchbenchmarkdatasetreliable] uses Reddit upvotes—confounding preference with popularity and timing—and covers only English. WritingBench [wu2025writingbench] spans six domains but mixes subjective creative tasks with objective functional ones (Academic & Engineering, Finance & Business). AlignBench [liu2023alignbench] evaluates Chinese LLM alignment but focuses on general capabilities rather than creative preference. WritingPreferenceBench advances beyond these by: (1) systematically neutralizing objective confounds through human validation, (2) providing cross-lingual coverage with consistent methodology, and (3) isolating purely subjective quality discrimination where prior benchmarks conflate multiple signals.
@@ 7 Conclusion
These results have theoretical and practical implications for preference learning. The 30 percentage point performance gap between architectures challenges the direct preference optimization paradigm [rafailov2023direct] and suggests that subjective domains require fundamentally different inductive biases than objective tasks. The failure of scale to improve performance (27B models underperform 7B variants) and the inability of reasoning-enhanced LLMs to surpass task-specific training indicate that current scaling laws may not apply to subjective preference modeling. Future work should investigate hybrid architectures combining the computational efficiency of discriminative models with the representational capacity of generative reasoning, and develop training objectives that explicitly encourage genre-invariant preference learning.